OpenStudy (anonymous):

Find the product of (4x + 3y)2.

OpenStudy (anonymous):

is that divided by 2?

OpenStudy (anonymous):

if not...it's just 8x + 6y

OpenStudy (anonymous):

\[(x+y)^2=x^2+2xy+y^2\]

OpenStudy (aroub):

16x^2 +24xy+9y^2

OpenStudy (anonymous):

(4x + 3y)2=8x+6y

OpenStudy (anonymous):

oh..raised to the 2nd power

OpenStudy (anonymous):

plz use ^2 to raise it up so we kknow